# How to Build a Quantum Supercomputer: Scaling Challenges and Opportunities

### v1

@@ III.5 Emulation of other FTQC protocols
Magic state preparation. A critical process in fault-tolerant quantum computation is preparation of high fidelity logical magic states. These states are produced by first employing a magic state preparation protocol Singh et al. (2022); Gidney (2023); Gidney et al. (2024), which produces relatively low fidelity logical magic states at small distances by employing physical $T$ gates. A large number of such logical states are then grown to higher distances and fed to magic state distillation units to prepare a lower number of higher fidelity magic states. These magic state distillation units are only able to perform if they are fed logical magic states of sufficient fidelity. It has been estimated that the 15-to-1 distillation units have acceptance probability
@@
where $P_{\text{magic}}$ and $P_{\text{Cliff}}$ are the logical error rates of input logical magic state and Clifford operations respectively Beverland et al. (2022). Hence, we need to produce logical magic states with error probability
@@
Whether magic states of this fidelity can be produced depends both on the specific magic state preparation protocol used and the hardware noise profile. We studied a protocol that cleverly exploits hook-injection errors to create high fidelity relatively low-distance magic states Gidney (2023) using Clifford simulations. This protocol first uses a physical $T$-gate to create a magic state on a small rotated surface code patch of distance $d_{1}$. It then post-selects the states for which no errors are detected and grows them to a larger distance $d$. The simulation results reported in Figure 21 show that the error rates increase with distance $d$, suggesting that the protocol is not fault tolerant, and explains why distillation units are needed instead of directly growing the magic states to a target distance. For simplicity we substitute the logical Clifford error rate with the logical error rate of the memory protocol in Equation 3 and draw the respective threshold curves for each of the hardware parameters studied in Figure 21. We observe that both target and $\Lambda_{18}$ parameter sets are significantly below their respective 15-to-1 distillation thresholds, while the baseline set is not.
@@
Figure 21: Performance of the magic state preparation protocols Gidney2023 Gidney (2023) and Gidney2024 Gidney et al. (2024) for the three parameter sets. Here, we fix $d_{1}=3$, because it yields the best performance for the baseline parameter set. The dashed curves indicate the distillation unit input threshold (3) for the baseline, target, and $\Lambda_{18}$ parameter set, respectively, where the difference is due to the fact that $P_{\text{Cliff}}$, estimated from the memory protocol logical error rate, depends on hardware noise. Observe that the baseline set is below the respective 15-to-1 distillation threshold only for the Gidney2024 protocol, while it is above threshold for Gidney2023.
@@
A more recent protocol Gidney et al. (2024) demonstrates significant improvements in the logical error rates for magic state preparation. This protocol introduces a number of improvements over past protocols, such as cleverly designed gradual growth stages and appropriate post-selections to ensure that error rates drop when increasing distances in practical regimes of error rates. We use the authors’ code to estimate the error rates for our three hardware parameter sets, also shown in Figure 21. We observe that all three hardware parameter sets are significantly below the 15-to-1 distillation threshold for this protocol and therefore the QREs in this paper use this protocol.
@@
Teleportation of logical qubits. To perform lattice surgery fault tolerantly, both space-like and time-like errors must be corrected. As discussed and numerically demonstrated in Chamberland and Campbell (2022), space-like errors are exponentially suppressed by increasing the code distance of the logical qubit. Similarly, time-like errors are exponentially reduced by increasing the number of stabilizater measurement rounds during the merge operation (see Figure 22, middle). In this context, the number of parity check cycles conducted while the two logical patches are merged is referred to as the temporal code distance. To evaluate the overall success rate of lattice surgery, we assess logical qubit teleportation under varying space–time parameters.
@@
Figure 22: Logical teleportation via lattice surgery for a rotated surface code of distance $d=3$ and bus width $b=1$. Illustrated are the states before merging (left), during merging (middle), and after splitting (right). The wire diagram of the teleporation quantum circuit including the relevant classical feedback and the needed recovery operations is also shown. We simulate the portion of the circuit before the dashed line using a circuit-level noise model. Note that at the dashed line invoking a decoder may flip the values $m_{1}$ and $m_{2}$. We do not simulate the circuity of the recovery operations $Z_{L}$ and $X_{L}$; instead, we measure all qubits at the dashed line to determine the performance of the teleporation.
@@
To determine the success rate of logical qubit teleportation, we estimate the average state infidelity defined by $P_{a}=\frac{1}{6}\sum_{\psi}P_{\psi}$, where $P_{\psi}$ represents the infidelity of teleporting each logical state $|\psi\rangle\in\{|0_{L}\rangle,|1_{L}\rangle,|+_{L}\rangle,|-_{L}\rangle,|+i_{L}\rangle,|-i_{L}\rangle\}$. Note that, the process fidelity is then given by $F_{p}=\frac{(D+1)(1-P_{a})-1}{D}=1-3/2P_{a}$, where $D=2$ is the dimension of the Hilbert space of the single qubit being teleported Nielsen (2002); Ryan-Anderson et al. (2024). Since accessing the $Y$ operator of the surface code is cumbersome Gidney (2024), we teleport the logical states $|\psi_{L}\rangle\in\{|0_{L}\rangle,|+_{L}\rangle\}$ using an $XX$ (rough) merge, as illustrated in Figure 22. From these simulations, we approximate the average state infidelity as $P_{a}\approx\frac{2}{3}(P_{+}+P_{0})$, providing an overestimate of the infidelity. The teleporation protocol we study is depicted in Figures 22 and 23 and outlined as follows:
@@
    Preparation: We begin by perfectly preparing the logical source state $|\psi_{L}\rangle\in\{|0_{L}\rangle,|+_{L}\rangle\}$ and the target state $|0_{L}\rangle$. These states are stabilized for $r_{pm}$ rounds, as show in in Figure 22 (left).
@@
    Merging: After the round $r_{pm}$, the bus data qubits are initialized in the physical $|0\rangle$ state and the rough edges of the two surfaces are merged. Then the entire surface is stabilized for $r_{m}$ rounds to determine (a possibly erroneous) measurement of the $X\otimes X$ observable with outcome $m_{1}$, as shown in Figure 22 (middle).
@@
    Splitting: After the round $r_{pm}+r_{m}$, the bus data qubits are measured in the $Z$ basis (splitting) and the remaining patches are stabilized for $r_{s}$ rounds. Note that this also results in a perfect round of syndrome measurements on the bus patches.
@@
    Projection: After the round $r_{pm}+r_{m}+r_{s}$, the source data qubits are measured in the $Z$-basis, to determine the (possibly erroneous) value of $m_{2}$. The projection of the source data qubits is illustrated in Figure 22 (right).
@@
    Recovery: At this point, by invoking a decoder, the values of $m_{1}$ and $m_{2}$ may be corrected and the recovery operations $Z_{L}$ and $X_{L}$ are conditionally applied.
@@
Figure 23: space–time diagrams for teleportation via lattice surgery. (a) space–time diagram of teleportation where $r_{s}=0$ QEC cycles are performed after splitting. The spatial dimensions are the same as shown in Figure 27, and the temporal dimension is divided into $r_{pm}$ pre-merge rounds and $r_{m}$ merge rounds. (b) The effect of buffer and decoder delays on lattice surgery fidelities. For each logical operation in the core processor, a teleportation is implemented involving a magic state (the left code patch) and data qubits (represented by the right code patch). The magic state may incur a delay $\tau_{b}$ in the buffer before the surgery is performed. The source (magic) state and the bus are measured out after the merge operation but target patches must await the decoder decisions, available after a decoder delay time, $\tau_{d}$. These idling patches must be protected using further stabilization rounds (quantum memory) during this period; therefore, accumulation of further errors is inevitable and must be taken into account by the compiler, FTQC emulator, and resource estimators.
@@
To estimate $P_{a}$ at large distances we simulate the above-mentioned teleportation protocol for varying spatial code distances $d$, $3\leq d\leq 15$, and temporal code distances $r_{m}$, $5\leq r_{m}\leq d$, fixing the bus width $b=3d$ and incorporating the circuit-level noise model detailed in Section III.2. We regress the following predictive models from the obtained numerical results of $P_{0}$ and $P_{+}$ values to predict the fidelity of the protocol at high distances Chamberland and Campbell (2022); Silva et al. (2024a):
$P_{0}$ $=\mu_{X}(2d+b)r_{m}\Lambda_{X}^{-(d+1)/2},$ (4)
$P_{+}$ $=\mu_{Z}d\Lambda_{Z}^{-(d+1)/2}+\mu_{T}db\Lambda_{T}^{-(r_{m}+1)/2}.$ (5)
@@
In our simulations the pre-merge stabilization rounds is fixed to $r_{pm}=1$, during which the two logical states are prepared (perfectly). We also use $r_{s}=0$ as illustrated Figure 23(a). For our benchmark purposes the recovery step is not performed, instead, the target data qubits are also measured in the basis corresponding to the initial source state $|\psi_{L}\rangle$ to determine the teleported logical state on the target.
@@
In Figure 24(a), we report the logical error rates of teleporting the states $|+_{L}\rangle$ and $|0_{L}\rangle$ (labeled $P_{+}$ and $P_{0}$, respectively) as a function of the code distance for $3\leq d\leq 15$ with corresponding bus width $b=3d$ and temporal code distance $r_{m}\in\{d,3d\}$. We highlight two observations from Figure 24(a).
@@
  • The error rates are suppressed as a function of code distance for both states. This indicates that the noise parameters are below threshold.
@@
  • Increasing $r_{m}$ decreases the teleportation fidelity of $|0_{L}\rangle$ while increasing the fidelity of teleporting $|+_{L}\rangle$ (see also Figure 24(d)).
@@
In Figure 24(b), we show the fitting for the $X$ and $Z$-type terms of Equations 4 and 5. This model predicts the $X$ and $Z$-type errors better at the high-$r_{m}$ regime (hence choosing the $r_{m}=3d$ data). Similarly, in Figure 24(c) we estimate the time-like error suppression term in Equation 5 at the high-distance regime, obtaining $\mu_{T}\approx 0.0273(6)$ and $\Lambda_{T}\approx 1.967(6)$. This information is sufficient to estimate the average error rate $P_{a}$ of high-distance teleportations.
@@
Figure 24: (a) Error suppression as a function of code distance $d$ for bus width $b=3d$, merged stabilization round $r_{m}\in\{d,3d\}$ for teleportation of $|0_{L}\rangle$ ($P_{0}$) and $|+_{L}\rangle$ ($P_{+}$). (b) An exponential fit for $P_{0}(r_{m}=3d)$ and $P_{+}(r_{m}=3d)$ using an exponential functions of the form $P_{0}\approx\mu_{X}d^{2}\Lambda_{X}^{-(d+1)/2}$ and $P_{+}\approx\mu_{Z}d\Lambda_{Z}^{-(d+1)/2}$. (c) An exponential suppression of $P_{+}$ as a function of the number of rounds for rounds $r_{m}<d$, for which we expect the logical error suppression of the form $P_{+}\approx\mu_{T}d^{2}\Lambda_{T}^{-(r_{m}+1)/2}$. (d) Logical teleportation error rates for teleporting the states $|+_{L}\rangle$ and $|0_{L}\rangle$ as a function of the merged stabilization rounds (temporal code distance) for the code of distance $d=7$ over a bus of width $b\in\{d,3d\}$ and the corresponding average state infidelities calculated using $\frac{2}{3}(P_{0}+P_{+})$.
@@
As another application, in Figure 24(d), we find the optimal number merge stabilization rounds $r_{m}$ for a given distance $d$. Here we choose $d=7$ and use two sizes for the bus $b\in\{d,3d\}$. We plot the estimated teleportation fidelity $P_{a}\approx\frac{2}{3}(P_{0}+P_{+})$ as a function of $r_{m}$ and note that the minimum of each curve is at $r_{m}>7$, highlighting the fact that the optimal number of QEC rounds for a lattice surgery operation with code distance $d$ may deviate from the commonly assumed value, $d$.
@@
For teleporation of magic states in our core processor, the source (magic) state has resided in the buffer for some average expected buffer delay time, $\tau_{b}$ (which can be as low as 1 clock cycle for balanced production and consumption of the magic states). The targets of teleporation are logical data qubit patches in the core processor for which further QEC rounds are executed until the decoder outcome is available. We denote this delay by $\tau_{d}$. Inclusion of the buffer and decoder delays and assuming an average rate for all types of surgeries results in the model
@@
which still distinguishes time-like and space-like errors but ignores the type of surgery; e.g., $XX$ merge or otherwise (Figure 23(b)).

### v2

@@ III.5 Emulation of other FTQC protocols
Magic state preparation. A critical process in FTQC is the preparation of high-fidelity logical magic states. These states are produced by first employing a magic state preparation protocol Singh et al. (2022); Gidney (2023); Gidney et al. (2024), which produces relatively low-fidelity logical magic states at small distances by employing physical $T$ gates. A large number of such logical states are then grown to greater distances and fed to magic state distillation units to prepare a lower number of higher-fidelity magic states. These magic state distillation units are able to perform only if they are fed logical magic states of sufficient fidelity. It has been estimated that the 15-to-1 distillation units have an acceptance probability
@@
where $P_{\text{magic}}$ and $P_{\text{Cliff}}$ are the logical error rates of input logical magic states and Clifford operations, respectively Beverland et al. (2022). Hence, we need to produce logical magic states with error probability
@@
Whether magic states of this fidelity can be produced depends both on the specific magic state preparation protocol used and the hardware noise profile. We studied a protocol that cleverly exploits hook-injection errors to create high-fidelity relatively low-distance magic states Gidney (2023) using Clifford simulations. This protocol first uses a physical $T$-gate to create a magic state on a small rotated surface code patch of distance $d_{1}$. It then post-selects the states for which no errors are detected and grows them to a larger code distance $d$. The simulation results reported in Figure 22 show that the error rates increase with $d$, suggesting that the protocol is not fault tolerant, and explains why distillation units are needed instead of directly growing the magic states to a target distance. For simplicity, we substitute the logical Clifford error rate with the logical error rate of the memory protocol in Equation 5 and draw the respective threshold curves for each of the hardware parameters shown in Figure 22. We observe that both the target and desired parameter sets are significantly below their respective 15-to-1 distillation thresholds, unlike the baseline set.
@@
Figure 22: Performance of the magic state preparation protocols Gidney2023 Gidney (2023) and Gidney2024 Gidney et al. (2024) for the three parameter sets. Here, we fix $d_{1}=3$, because it yields the best performance for the baseline parameter set. The dashed curves indicate the distillation unit input threshold (5) for the baseline, target, and desired parameter set, respectively, where the difference is due to the fact that $P_{\text{Cliff}}$, estimated from the memory protocol logical error rate, depends on hardware noise. We observe that the baseline set is below the respective 15-to-1 distillation threshold only for the Gidney2024 protocol; it is above the threshold for the Gidney2023 protocol.
@@
A more recent protocol Gidney et al. (2024) demonstrates significant improvements in the logical error rates for magic state preparation. This protocol introduces a number of improvements over past protocols, such as cleverly designed gradual growth stages and appropriate post-selections to ensure that error rates drop when increasing distances in practical regimes of error rates. We use the authors’ code to estimate the error rates for our three hardware parameter sets, also shown in Figure 22. We observe that all three parameter sets are significantly below the 15-to-1 distillation threshold for this protocol and for this reason the QREs in this paper use this protocol.
@@
Teleportation of logical qubits. To perform lattice surgery fault tolerantly, both space-like and time-like errors must be corrected. As discussed and numerically demonstrated in Ref. Chamberland and Campbell (2022), space-like errors are exponentially suppressed by increasing the code distance of the logical qubits. Similarly, time-like errors are exponentially reduced by increasing the number of stabilizer measurement rounds during the merge operation (see Figure 23). In this context, the number of parity-check cycles conducted while the two logical patches are merged is referred to as the temporal code distance. To evaluate the overall success rate of lattice surgery, we assess logical qubit teleportation under varying space–time parameters.
@@
Figure 23: Logical teleportation via lattice surgery for a rotated surface code of distance $d=3$ and bus width $b=1$. Illustrated are the states prior to merging (left), during merging (middle), and after splitting (right). The wire diagram of the teleportation quantum circuit, including the relevant classical feedback and the needed recovery operations, is also shown. We simulate the portion of the circuit before the dashed line using a circuit-level noise model. Note that, at the dashed line, invoking a decoder may flip the values of $m_{1}$ and $m_{2}$. We do not simulate the circuity of the recovery operations $Z_{L}$ and $X_{L}$; instead, we measure all qubits at the dashed line to determine the performance of the teleportation.
@@
To determine the success rate of logical qubit teleportation, we estimate the average state infidelity defined by $P_{a}=\frac{1}{6}\sum_{\psi}P_{\psi}$, where $P_{\psi}$ represents the infidelity of teleporting each logical state $|\psi\rangle\in$$\{|0_{L}\rangle,|1_{L}\rangle,|+_{L}\rangle,|-_{L}\rangle,|+i_{L}\rangle,|-i_{L}\rangle\}$. Note that the process fidelity is then given by $F_{p}=\frac{(D+1)(1-P_{a})-1}{D}=1-3/2P_{a}$, where $D=2$ is the dimension of the Hilbert space of the single qubit being teleported Nielsen (2002); Ryan-Anderson et al. (2024). Since accessing the $Y$ operator of the surface code is cumbersome Gidney (2024), we teleport the logical states $|\psi_{L}\rangle\in\{|0_{L}\rangle,|+_{L}\rangle\}$ using an $XX$ (rough) merge, as illustrated in Figure 23. From these simulations, we approximate the average state infidelity as $P_{a}\approx\frac{2}{3}(P_{+}+P_{0})$, providing an overestimate of the infidelity. The teleportation protocol we study is illustrated in Figures 23 and 24(a) and outlined as follows:
@@
    Preparation (Figure 23, left side): We begin by preparing the logical source state $|\psi_{L}\rangle\in\{|0_{L}\rangle,|+_{L}\rangle\}$ and the target patch in the state $|0_{L}\rangle$. In this pre-merging step, these states are stabilized for $r_{\text{pm}}$ rounds.
@@
    Merging (Figure 23, middle): After the $r_{\text{{pm}}}$ rounds, the bus data qubits are respectively initialized in the physical $|0\rangle$ state, and the rough edges of the two surfaces are merged. Then, the entire surface is stabilized for $r_{\text{m}}$ rounds to determine (a possibly erroneous) measurement of the $X\otimes X$ observable with an outcome $m_{1}$.
@@
    Splitting (Figure 23, right side): After the $r_{\text{pm}}+r_{\text{m}}$ rounds, the bus data qubits are respectively measured in the $Z$ basis (splitting) and the remaining patches are stabilized for $r_{\text{s}}$ rounds. Note that this also results in a perfect round of syndrome measurements on the bus patches.
@@
    Projection (Figure 23, right side): After the $r_{\text{pm}}+r_{\text{m}}+r_{\text{s}}$ rounds, the source data qubits are respectively measured in the $Z$-basis, to determine the (possibly erroneous) value of an outcome $m_{2}$.
@@
    Recovery (Figure 23, right side): At this point, by invoking a decoder, the values of $m_{1}$ and $m_{2}$ may be corrected, and the recovery operations $Z_{L}$ and $X_{L}$ are conditionally applied.
@@
Figure 24: Space–time diagrams of teleportation via lattice surgery. (a) Space–time diagram of teleportation where $r_{\text{s}}=0$ QEC cycles are performed after splitting. The spatial dimensions are the same as shown in Figure 28, and the temporal dimension is divided into $r_{\text{pm}}$ pre-merge rounds and $r_{\text{m}}$ merge rounds. (b) Effect of buffer and decoder delays on lattice surgery fidelities. For each logical operation in the core processor, a teleportation is implemented involving a magic state (represented by the left code patch) and data qubits (represented by the right code patch). The magic state may incur a delay $\tau_{b}$ in the buffer before the surgery is performed. The magic state and the bus are measured after the merge operation, but target patches must await the decoder decisions, available after a decoder delay time $\tau_{d}$. These idling patches must be protected by a quantum memory protocol using further stabilization rounds during this period; therefore, the accumulation of further errors is inevitable and must be taken into account by the compiler, FTQC emulator, and resource estimators.
@@
To estimate $P_{a}$ at large distances, we simulate the mentioned teleportation protocol for varying spatial code distances $d$, temporal code distances $r_{\text{m}}$, and bus widths $b$. We then regress the following predictive models from the obtained numerical results of $P_{0}$ and $P_{+}$ values to predict the fidelity of the protocol at large distances Chamberland and Campbell (2022); Silva et al. (2024a):
$P_{0}$ $=\mu_{X}(2d+b)r_{\text{m}}\Lambda_{X}^{-(d+1)/2},$ (6)
$P_{+}$ $=\mu_{Z}d\Lambda_{Z}^{-(d+1)/2}+\mu_{T}db\Lambda_{T}^{-(r_{\text{m}}+1)/2}.$ (7)
@@
In our simulations, the number of pre-merge stabilization rounds is fixed at $r_{\text{pm}}=1$, during which the two logical states are prepared. For the remaining rounds, we incorporate the circuit-level noise model detailed in Appendix E using the baseline noise parameters. We also set $r_{\text{s}}=0$. For our benchmarking purposes, the recovery step is not performed; instead, the target data qubits are also measured in the basis corresponding to the initial source state $|\psi_{L}\rangle$ to determine the teleported logical state at the target patch.
@@
Figure 25(a) shows the logical error rates in teleporting the states $|+_{L}\rangle$ and $|0_{L}\rangle$ (labeled $P_{+}$ and $P_{0}$, respectively) as a function of the code distance for $3\leq d\leq 15$, with a corresponding bus width $b=3d$ and a temporal distance $r_{\text{m}}\in\{d,3d\}$. We highlight two observations from Figure 25(a):
@@
  • The error rates are suppressed as a function of code distance for both states. This indicates that the noise parameters are below the threshold.
@@
  • Increasing $r_{\text{m}}$ decreases the teleportation fidelity of $|0_{L}\rangle$ while increasing the fidelity of teleporting $|+_{L}\rangle$ (see also Figure 25(d)).
@@
Figure 25(b) shows the fitting for the $X$- and $Z$-type terms of Equations 6 and 7. This model predicts the $X$- and $Z$-type errors better in the high-$r_{\text{m}}$ regime (hence our choice of the $r_{\text{m}}=3d$ data). Similarly, Figure 25(c) shows an estimate of the time-like error suppression term in Equation 7 in the high-distance regime, yielding the values $\mu_{T}\approx 0.0273(6)$ and $\Lambda_{T}\approx 1.967(6)$. This information is sufficient to estimate the average error rate $P_{a}$ of high-distance teleportations.
@@
Figure 25: (a) Error suppression as a function of code distance $d$ for a bus width $b=3d$ and for $r_{\text{m}}\in\{d,3d\}$ merge rounds, for the teleportation of $|0_{L}\rangle$ ($P_{0}$) and $|+_{L}\rangle$ ($P_{+}$). (b) Exponential fit for $P_{0}(r_{\text{m}}=3d)$ and $P_{+}(r_{\text{m}}=3d)$ using exponential functions of the form $P_{0}\approx\mu_{X}d^{2}\Lambda_{X}^{-(d+1)/2}$ and $P_{+}\approx\mu_{Z}d\Lambda_{Z}^{-(d+1)/2}$. (c) Exponential suppression of $P_{+}$ as a function of the number of rounds for $r_{\text{m}}<d$, for which we expect the logical error suppression of the form $P_{+}\approx\mu_{T}d^{2}\Lambda_{T}^{-(r_{\text{m}}+1)/2}$. (d) Logical teleportation error rates for teleporting the states $|0_{L}\rangle$ and $|+_{L}\rangle$ as a function of the merge stabilization rounds (temporal distance) for a code distance $d=7$ over a bus width $b\in\{d,3d\}$ and the corresponding average state infidelities calculated using $\frac{2}{3}(P_{0}+P_{+})$.
@@
As another application,Figure 25(d) shows the optimal number of merge stabilization rounds $r_{\text{m}}$ for a given distance $d$. Here, we choose $d=7$ and consider two sizes for the bus $b\in\{d,3d\}$. We plot the estimated teleportation fidelity $P_{a}\approx\frac{2}{3}(P_{0}+P_{+})$ as a function of $r_{\text{m}}$ and note that the minimum of each curve is at values of $r_{\text{m}}>7$, highlighting the fact that the optimal number of QEC rounds for a lattice surgery operation with code distance $d$ may actually deviate from the commonly assumed value $d$.
@@
For the teleportation of magic states in the core processor, the magic state has resided in the buffer for some average expected buffer delay time $\tau_{b}$, which can be as low as 1 clock cycle for balanced production and consumption of the magic states. The targets of teleportation are logical data qubit patches in the core processor for which further QEC rounds are executed until the decoder outcome becomes available. We denote this delay by $\tau_{d}$. Inclusion of the buffer and decoder delays and assuming an average rate for all types of surgeries results in the model
@@
which still distinguishes time-like and space-like errors but ignores the type of surgery (see Figure 24(b)).